# Traces of vanishing Hölder spaces

### v2

@@ Abstract.
[…]
@@ 1.1. Introduction
The vanishing Hölder spaces $\dot{\operatorname{VC}}_{\Gamma}^{\omega}(X,Y),$ for $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ for normed spaces $X,Y$ and for many moduli of continuity $\omega,$ were recently defined and studied by the two last named authors [MudOik24]. In particular, these vanishing Hölder spaces were shown to often provide a complete description of those Hölder functions approximable by Lipschitz or smooth or boundedly supported functions. In this article we continue to study the vanishing spaces $\dot{\operatorname{VC}}_{\Gamma}^{\omega}(X,Y),$ when $X=\mathbb{R}^{n}$ but $Y$ is allowed to be an arbitrary normed space.
@@
One perspective to our results is as follows. The scale of Hölder spaces $\dot{C}^{0,\omega}$ can be seen to measure the smoothness of a function relative to any fixed modulus $\omega(t)>0,$ and when $\omega(t)=0$ the space $\mathrm{BMO}$ of bounded mean oscillations is often substituted as the zeroth order smoothness measure. As opposed to real-valued Hölder functions $\dot{C}^{0,\omega}(E),$ with $E\subset\mathbb{R}^{n}$ arbitrary, being extendable to the whole ambient space $\dot{C}^{0,\omega}(\mathbb{R}^{n}),$ for example by the classical infimal convolution formula $x\mapsto\inf_{y\in E}\{f(y)+M\omega(|x-y|)\},$ this is far from being true for functions in $\mathrm{BMO}.$ Indeed, given a connected open set (a domain) $\Omega\subset\mathbb{R}^{n},$ a classical theorem of Jones [Jon80Ext] tells us that a bounded linear extension operator $L:\mathrm{BMO}(\Omega)\to\mathrm{BMO}(\mathbb{R}^{n})$ exists if and only if $\Omega$ is uniform. With $\mathrm{BMO},$ geometry of the domain appears naturally. A recent result of Dalia and Gafni [ButDaf21, Theorem 3.] (see also [ButDaf23]) extends Jone’s result by providing a single bounded extension operator on $\mathrm{BMO}(\Omega)$ and also on several of its subspaces determined by approximability by nice functions, or by having appropriate vanishing mean oscillations. Our results are in the same spirit as Butaev’s and Gafni’s, but opposed to Jones’; we consider vanishing subspaces, but get rid of all geometry by reintroducing an order of smoothness $\omega(t)>0.$
@@
It follows from our results that a single linear and bounded extension operator exists for the intersection $\dot{\operatorname{VC}}^{\omega}$ of the three vanishing scales. In particular, for the Hölder moduli $\alpha(t):=t^{\alpha},$ for $\alpha\in(0,1),$ [MudOik24, Theorem 1.13.] states that $\dot{\operatorname{VC}}^{\alpha}(\mathbb{R}^{n})=\operatorname{VMO}^{\alpha}(\mathbb{R}^{n}),$ [^1] We remark that notation-wise $\operatorname{VMO}^{\alpha}=\operatorname{CMO}^{\alpha}$ for Guo et al. where the latter fractional vanishing mean oscillation space was defined in Guo et al. [GHWY21, Theorem 1.7.] as the space that characterizes the off-diagonal $L^{p}\to L^{q}$ (for $1<p<q<\infty$) compactness of commutators of singular integrals. The study of these vanishing Hölder scales is thus also motivated by the needs of singular integrals and harmonic analysis on $\mathbb{R}^{n}.$ Moreover, the classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are fundamental in the study of Lipschitz algebras in metric spaces; we refer to the monographs by Weaver [Weaverbook, Chapters 4 and 8], and for an exposition of these little Hölder spaces in the setting of Ahlfors regular sets we refer to D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3].
@@
For the extension of $\dot{C}^{0,\omega}$ mappings between subsets of Hilbert spaces, preserving the $\dot{C}^{0,\omega}$-seminorms, see Grünbaum and Zarantonello [GruZar68], and the monograph [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss. Concerning classical (and fundamental) results on $C^{m,\omega}$ extension of jets in $\mathbb{R}^{n},$ we refer to the work of Whitney [Wh34] and Glaeser [Gl58]. For Whitney-type extensions of jets generated by Sobolev functions in $\mathbb{R}^{n}$, see, for instance, the recent paper of Shvartsman [Shv17]. For infinite dimensional results on $C^{1,\omega}$ extension of jets, see the recent paper [AzaMud21] by Azagra and the second named author of the present paper. For extension results for functions (instead of jets) of order $C^{m}$ or $C^{m,\omega}$, or for Sobolev functions, see for instance the papers by Brudnyi and Shvartsman [BruShv01], by Fefferman [Fef05, Fef06], by Fefferman, Israel and Luli [FefIsrLul14].
@@ 1.2. Basic definitions and main results
It is clear that $\|\cdot\|_{\dot{\operatorname{J}}^{m,\omega}(E,V)}$ defines a seminorm on $\dot{\operatorname{J}}^{m,\omega}(E,V)$, and it can be made an actual norm if we fix a point $x_{0}\in E$ and define
@@
In the case $m=0,$ the trace space $\dot{\operatorname{J}}^{0,\omega}(E,V)$ coincides with the homogeneous Hölder space $\dot{C}^{0,\omega}(E,V)$, consisting of those $f:E\to V$ such that
@@
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{small}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}x\not=y\in E\\ |x-y|\leq\delta\end{subarray}}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=0\Big{\}},$
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{large}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}x\not=y\in E\\ |x-y|\geq\delta\end{subarray}}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-z|)}=0\Big{\}},$
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{far}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to\infty}\sup_{\min(|x|,|y|)>\delta}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=0\Big{\}}.$
@@
Given an $m$-jet $\{A_{k}:E\to\mathcal{L}^{k}(\mathbb{R}^{n},V)\}_{k=0}^{m}$ on $E$, what necessary and sufficient conditions guarantee the existence of $F\in\dot{\operatorname{VC}}^{m,\omega}_{\Gamma}(\mathbb{R}^{n},V)$ so that $D^{k}F$ restricted to $E$ agrees with $A_{k},$ for each $k=0,\ldots,m$? Does there exist an extension operator that works simultaneously for all of the three vanishing subspaces? Can such an extension be defined by means of a linear operator?
@@
$\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V):=\Big{\{}\{A_{k}\}_{k=0}^{m}\in\dot{\operatorname{J}}^{m,\omega}(E,V)\,:\,\lim_{\delta\to\infty}\sup_{\min(|x|,|y|)\geq\delta}\frac{R\left(\{A_{k}\}_{k=0}^{m},x,y\right)}{\omega(|x-y|)}=0\Big{\}}.$
@@
The following lemma will be used both in this and the next section.
@@
Let $\mathcal{A}=\{A_{k}\}_{k=0}^{m}\in\dot{\operatorname{VJ}}^{m,\omega}(E,V)$ and $\varepsilon>0.$ Let $M=M(\varepsilon)$ be such that if $u,v\in E$ and $|u|,|v|>M,$ then
$R\left(\mathcal{A},u,v\right)<\varepsilon\omega(|u-v|).$
Now we consider arbitrary $x,y\in E,$ let $K\gg M$ be a large parameter that we will determine later, and let us assume
$\max(|x|,|y|)\geq K\gg M\geq\min(|x|,|y|).$ (6)
Notice also that the condition $\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V)$ is not vaguous if and only if $E$ is unbounded. So for each $M>0,$ there is $z\in E\setminus B(0,M)$ (independent of $x$ and $y$). Assume further that $K\gg|z|$ from now on. Bearing in mind that $x,y$ satisfy (6), we start by checking the case $|x|\geq K,$ and $|y|\leq M$ first. Because of the properties of $K,$ $M,$ $x,y,$ and $z,$ it is clear that
$|y-z|\leq|x-y|\quad\text{and}\quad C^{-1}|x-y|\leq|x-z|\leq C|x-y|;$ (7)
for some absolute constant $C.$ Now, using that $x,z\in E\setminus B(0,M)$, the relations (7) and Remark 2.1, we have, for every $0\leq k\leq m,$
@@
$\leq\Big{\|}A_{k}(x)-\sum_{j=0}^{m-k}\frac{1}{j!}A_{k+j}(z)(x-z)^{j}\Big{\|}+\Big{\|}\sum_{j=0}^{m-k}\frac{1}{j!}A_{k+j}(z)\left((y-z)^{j}-(x-z)^{j}\right)\Big{\|}$
$\leq\varepsilon|x-z|^{m-k}\,\omega(|x-z|)+C(m)\max_{0\leq j\leq m-k}\sup_{B(0,|z|)}\|A_{k+j}\|\left(|y-z|^{j}+|x-z|^{j}\right),$
$\lesssim_{C_{\omega},m,E,\|\mathcal{A}\|_{\dot{\operatorname{VJ}}^{m,\omega}(E,V)}}\varepsilon|x-y|^{m-k}\omega(|x-y|)+C(|z|,\mathcal{A})|x-y|^{m-k}.$ (8)
Since $K$ was so that $K\gg|z|$, we assume that $K$ is large to guarantee $C(|z|,\mathcal{A})\leq\varepsilon\omega(K).$ Thus $C(|z|,\mathcal{A})\lesssim_{C_{\omega}}\varepsilon\omega(|x-y|),$ and from (2), we derive the bound
@@
Dividing by the term $|x-y|^{m-k}\omega(|x-y|)$, and taking the maximum among $k=0,\ldots,m,$ we arrive at $R(\mathcal{A},x,y)/\omega(|x-y|)\lesssim\varepsilon.$
Let us now study the (non-symmetric) case $|y|\geq K,$ and $|x|\leq M.$ Let us write
@@
$\lesssim_{m}\sum_{j=0}^{m-k}\Big{\|}A_{k+j}(y)-\sum_{l=0}^{m-k-j}\frac{1}{l!}A_{k+j+l}(x)(y-x)^{l}\Big{\|}|x-y|^{j}+\max_{1\leq l\leq m}\sup_{|u|\leq M}\|A_{l}(u)\||x-y|^{m-k}.$ (10)
For the first term in (2), we use the estimates (9) for each $j=0,\ldots,m-k$, but swapping the roles of $x$ and $y$ (note that the estimates of (9) were obtained for $|x|\geq K$ and $|y|\leq M$). We thus obtain, for $K$ large enough that
@@
As for the second term in (2), recall that $K\gg M$ and so, bearing in mind property (5) and Remark 2.1, we can say that $\max_{1\leq l\leq m}\sup_{|u|\leq M}\|A_{l}(u)\|\leq\varepsilon\omega(K)\lesssim_{C_{\omega}}\varepsilon\omega(|x-y|).$ Dividing by $|x-y|^{m-k}\omega(|x-y|)$ and combining the above two cases, we conclude that
@@
So, let us assume that $F\in\dot{\operatorname{VC}}^{m,\omega}_{\Gamma}(\mathbb{R}^{n},V)$ and prove that the restriction of $F$ to $E$ satisfies the properties from Definition 1.6. While the proof in the case $\Gamma=\operatorname{small}$ is immediate from Taylor’s theorem, for the scales $\Gamma=\operatorname{large}$ and $\Gamma=\operatorname{far}$, we need to study a couple of subcases separately.
@@
Let $x,y\in\mathbb{R}^{n}$ are such that $|x|,|y|\geq M$, for certain $M\gg R$ that we will determine later. Let $z\in[x,y]$ be as on the right-hand side of (11). In the case where $\max(|y|,|z|)\geq R,$ then

### v3

@@ Abstract.
For an arbitrary subset $E\subset\mathbb{R}^{n},$ we introduce and study the three vanishing subspaces of the Hölder space $\dot{C}^{0,\omega}(E)$ consisting of those functions for which the ratio $|f(x)-f(y)|/\omega(|x-y|)$ vanishes, when $(1)$ $|x-y|\to 0$ , $(2)$ $|x-y|\to\infty$ or $(3)$ $\min(|x|,|y|)\to\infty.$ We prove that the Whitney extension operator maps each of these vanishing subspaces from $E$ to the corresponding vanishing spaces defined on the whole ambient space $\mathbb{R}^{n}.$ In fact, this follows as the zeroth order special case of a more general problem involving higher order derivatives. As a consequence, we obtain complete characterizations of approximability of Hölder functions $\dot{C}^{0,\omega}(E)$ by Lipschitz and boundedly supported functions.
@@ 1.1. Introduction
The vanishing Hölder spaces $\dot{\operatorname{VC}}_{\Gamma}^{\omega}(X,Y),$ for the scales $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ for normed spaces $X,Y$ and for many moduli of continuity $\omega,$ were recently defined and studied by the two last named authors [MudOik24]. In particular, these vanishing Hölder spaces were shown to often provide a complete description of those Hölder functions approximable by Lipschitz or smooth or boundedly supported functions. In this article we continue to study the vanishing Hölder spaces $\dot{\operatorname{VC}}_{\Gamma}^{\omega}(X,Y),$ when $X=\mathbb{R}^{n}$ but $Y$ is allowed to be an arbitrary normed space.
@@
One perspective to our results is as follows. The scale of Hölder spaces $\dot{C}^{0,\omega}$ can be seen to measure the smoothness of a function relative to the modulus $\omega(t)>0,$ and when $\omega(t)=0$ the space $\mathrm{BMO}$ of bounded mean oscillations is often substituted as the zeroth order way of measuring smoothness. As opposed to real-valued Hölder functions $\dot{C}^{0,\omega}(E,\mathbb{R}^{n}),$ with $E\subset\mathbb{R}^{n}$ arbitrary, being extendable to the whole ambient space $\dot{C}^{0,\omega}(\mathbb{R}^{n},\mathbb{R})$ by the classical infimal convolution formula $x\mapsto\inf_{y\in E}\{f(y)+M\omega(|x-y|)\},$ for example, this is far from being true for functions in $\mathrm{BMO}.$ Indeed, given a connected open set (a domain) $\Omega\subset\mathbb{R}^{n},$ a classical theorem of Jones [Jon80Ext] tells us that a bounded linear extension operator $L:\mathrm{BMO}(\Omega)\to\mathrm{BMO}(\mathbb{R}^{n})$ exists if and only if $\Omega$ is uniform. With $\mathrm{BMO},$ geometry of the domain appears naturally. A recent result of Butaev and Gafni [ButDaf21, Theorem 3] (see also [ButDaf23]) extends Jones’ result by providing a single bounded extension operator on $\mathrm{BMO}(\Omega)$ and also on several of its subspaces determined by approximability by nice functions, or by having appropriate vanishing mean oscillations. Our results are in the same spirit as Butaev’s and Gafni’s, but opposed to Jones’; we consider vanishing subspaces, but get rid of all geometry by reintroducing an order of smoothness $\omega(t)>0.$
@@
It follows from our results that a single linear and bounded extension operator exists for the intersection $\dot{\operatorname{VC}}^{\omega}$ of the three vanishing scales. In particular, for the Hölder moduli $\alpha(t):=t^{\alpha},$ for $\alpha\in(0,1),$ [MudOik24, Theorem 1.13] states that $\dot{\operatorname{VC}}^{\alpha}(\mathbb{R}^{n})=\operatorname{VMO}^{\alpha}(\mathbb{R}^{n}),$ [^1] We remark that notationally $\operatorname{VMO}^{\alpha}=\operatorname{CMO}^{\alpha}$ for Guo et al. where the latter fractional vanishing mean oscillation space was defined in Guo et al. [GHWY21, Theorem 1.7] and shown to characterizes the off-diagonal $L^{p}\to L^{q}$ (for $1<p<q<\infty$) compactness of commutators of many singular integral operators. We also point out that the identification $\dot{C}^{0,\alpha}(\mathbb{R}^{n},\mathbb{R})=\operatorname{BMO}^{\alpha}(\mathbb{R}^{n},\mathbb{R})$ of the Hölder spaces with the spaces of fractional bounded mean oscillation is due to N. G. Meyers [Mey1964]. Moreover, the classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are fundamental in the study of Lipschitz algebras in metric spaces; we refer to the monographs by Weaver [Weaverbook, Chapters 4 and 8], and for an exposition of these little Hölder spaces in the setting of Ahlfors regular sets we refer to D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3].
@@
For the extension of $\dot{C}^{0,\omega}$ mappings between subsets of Hilbert spaces, preserving the $\dot{C}^{0,\omega}$-seminorms, see Grünbaum and Zarantonello [GruZar68], and the monograph [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss. The fundamental starting point for $C^{m}$ extension of jets in $\mathbb{R}^{n}$ is the celebrated Whitney Extension Theorem [Wh34]. For Whitney-type $C^{m,\omega}$ extensions of jets in $\mathbb{R}^{n},$ we refer to the work of Glaeser [Gl58], and for Whitney-type extensions of jets generated by Sobolev functions in $\mathbb{R}^{n}$, see, for instance, the recent paper of Shvartsman [Shv17]. For infinite dimensional results on $C^{1,\omega}$ extension of jets, see the recent paper [AzaMud21] by Azagra and the second named author of the present paper. For extension results for functions (instead of jets) of order $C^{m}$ or $C^{m,\omega}$, or for Sobolev functions, see for instance the papers by Brudnyi and Shvartsman [BruShv01], by Fefferman [Fef05, Fef06], by Fefferman, Israel and Luli [FefIsrLul14].
@@ Acknowledgements
We are grateful to the anonymous referee for carefully reading the manuscript and for their comments, which led to improvements in both clarity and presentation.
@@ 1.2. Basic definitions and main results
It is convenient to clarify that each term $A_{k+j}(y)(x-y)^{j}$ in the sum of (1) is understood as the symmetric $k$-linear mapping $(\mathbb{R}^{n})^{k}\to V$ given by:
$A_{k+j}(y)(x-y)^{j}(u_{1},\ldots,u_{k}):=A_{k+j}(y)\Big{(}\overbrace{x-y,\ldots,x-y}^{j\text{ times}},u_{1},\ldots,u_{k}\Big{)},\quad(u_{1},\ldots,u_{k})\in(\mathbb{R}^{n})^{k}.$
In particular, the comparison of the sum of these mappings with $A_{k}(x)$ in (1) is well-defined. In addition, since $A_{k+j}(y)$ is a symmetric $(k+j)$-linear mapping $(\mathbb{R}^{n})^{k+j}\to V$, one can distribute the $j$-many vectors $x-y$ among the $k+j$ entries in any order, resulting in the same symmetric $k$-linear mapping as above.
Moreover, it is clear that $\|\cdot\|_{\dot{\operatorname{J}}^{m,\omega}(E,V)}$ defines a seminorm on $\dot{\operatorname{J}}^{m,\omega}(E,V)$, and it can be made an actual norm if we fix a point $x_{0}\in E$ and define
@@
In the case $m=0,$ the trace space $\dot{\operatorname{J}}^{0,\omega}(E,V)$ coincides with the homogeneous Hölder space $\dot{C}^{0,\omega}(E,V)$, consisting of those functions $f:E\to V$ such that
@@
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{small}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}x\not=y\in\mathbb{R}^{n}\\ |x-y|\leq\delta\end{subarray}}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=0\Big{\}},$
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{large}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}x,y\in\mathbb{R}^{n}\\ |x-y|\geq\delta\end{subarray}}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=0\Big{\}},$
$\dot{\operatorname{VC}}^{m,\omega}_{\operatorname{far}}(\mathbb{R}^{n},V):=\Big{\{}F\in\dot{C}^{m,\omega}(\mathbb{R}^{n},V)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}x\not=y\in\mathbb{R}^{n}\\ \min(|x|,|y|)\geq\delta\end{subarray}}\frac{\|D^{m}F(x)-D^{m}F(y)\|}{\omega(|x-y|)}=0\Big{\}}.$
@@
Given an $m$-jet $\{A_{k}:E\to\mathcal{L}^{k}(\mathbb{R}^{n},V)\}_{k=0}^{m}$ on $E$, a modulus of continuity $\omega$, and one of the scales $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ defined above, what necessary and sufficient conditions guarantee the existence of $F\in\dot{\operatorname{VC}}^{m,\omega}_{\Gamma}(\mathbb{R}^{n},V)$ so that $D^{k}F$ restricted to $E$ agrees with $A_{k},$ for each $k=0,\ldots,m$? Does there exist an extension operator that works simultaneously for all of the three vanishing subspaces? Can such an extension be defined by means of a linear operator?
@@
$\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V):=\Big{\{}\{A_{k}\}_{k=0}^{m}\in\dot{\operatorname{J}}^{m,\omega}(E,V)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}x\not=y\in E\\ \min(|x|,|y|)\geq\delta\end{subarray}}\frac{R\left(\{A_{k}\}_{k=0}^{m},x,y\right)}{\omega(|x-y|)}=0\Big{\}}.$
@@
The following Lemma 2.2 is extremely useful both in this and the next section. It shows that the minimum $\min(|x|,|y|)\geq\delta$ defining the vanishing condition $\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V)$ can be replaced with the maximum $\max(|x|,|y|)\geq\delta.$ We also note that the proof in the case $m=0$ is simpler due to the symmetry of the corresponding $R(\mathcal{A},x,y).$ For a proof in this special case with $E=\mathbb{R}^{n},$ we direct the interested reader to [MudOik24, Lemma 2.2.].
@@
First notice that the vanishing condition of $\dot{\operatorname{VJ}}_{\operatorname{far}}^{m,\omega}(E,V)$ is not vacuous if and only if $E$ is unbounded, and so $E$ is assumed to be unbounded in this proof. Let $\mathcal{A}=\{A_{k}\}_{k=0}^{m}\in\dot{\operatorname{VJ}}^{m,\omega}(E,V)$ and $\varepsilon>0.$ Let $M=M(\varepsilon)$ be such that if $u,v\in E$ and $|u|,|v|>M,$ then
$R\left(\mathcal{A},u,v\right)<\varepsilon\omega(|u-v|).$ (6)
Since $E$ is unbounded, we can fix a point $z_{0}\in E\setminus B(0,M)$. Now, by Remark 2.1 and the assumption (5), we can select $K>0$ large enough (and depending on $\varepsilon,M,\omega,E,\mathcal{A}$) so that
$K\geq 4\max\left(M,|z_{0}|,1\right)\quad\text{and}\quad\max_{0\leq j\leq m}\sup_{z\in E,\,|z|\leq|z_{0}|}\|A_{j}(z)\|\leq\varepsilon\omega(K).$ (7)
Because $|z_{0}|\geq M,$ the last bound is also true when $|z_{0}|$ is replaced by $M.$ (Morally speaking, we take $K\gg M+|z_{0}|$ to be much bigger than both $M$ and $|z_{0}|.$)
Now we consider arbitrary $x,y\in E$ with the lower bound $\max(|x|,|y|)\geq K.$ Notice that, for our purposes, we can also assume that $M\geq\min(|x|,|y|),$ as otherwise the desired estimate is true already by (6). So, for future reference, we are assuming that:
$\max(|x|,|y|)\geq K\quad\text{and}\quad M\geq\min(|x|,|y|).$ (8)
Notice that $R(\mathcal{A},x,y)$ is not symmetric with respect to $(x,y),$ and so we must study two non-symmetric cases.
We start by checking the case $|x|\geq K$ and $|y|\leq M$ first. Using the first inequality of (7), the fact $|z_{0}|\geq M$, and the triangle inequality, we see that
$|y-z_{0}|\leq|x-y|\quad\text{and}\quad C^{-1}|x-y|\leq|x-z_{0}|\leq C|x-y|;$ (9)
for some absolute constant $C>0.$ Now, using that $x,z_{0}\in E\setminus B(0,M)$, the relations (9) and the choice of $K$ in (7), we have, for every $0\leq k\leq m,$
@@
$\leq\Big{\|}A_{k}(x)-\sum_{j=0}^{m-k}\frac{1}{j!}A_{k+j}(z_{0})(x-z_{0})^{j}\Big{\|}+\Big{\|}\sum_{j=0}^{m-k}\frac{1}{j!}A_{k+j}(z_{0})\left((y-z_{0})^{j}-(x-z_{0})^{j}\right)\Big{\|}$
$\leq\varepsilon|x-z_{0}|^{m-k}\,\omega(|x-z_{0}|)+C(m)\max_{0\leq j\leq m-k}\sup_{z\in E,\,|z|\leq|z_{0}|}\|A_{k+j}(z)\|\left(|y-z_{0}|^{j}+|x-z_{0}|^{j}\right),$
$\lesssim_{C_{\omega},m,E,\|\mathcal{A}\|_{\dot{\operatorname{VJ}}^{m,\omega}(E,V)}}\varepsilon|x-y|^{m-k}\omega(|x-y|)+\varepsilon\omega(K)\max_{0\leq j\leq m-k}|x-y|^{j}.$ (10)
Since $|x-y|\geq K-M\geq K/2\geq 1$ by (7), we have that $w(K)\lesssim_{\omega}\omega(|x-y|)$ (see condition (3) of $\omega$) and also $|x-y|^{j}\leq|x-y|^{m-k}$ for all $0\leq j\leq m-k.$ These applied to (2) yield the bound
@@
Dividing by the term $|x-y|^{m-k}\omega(|x-y|)$, and taking the maximum among $k=0,\ldots,m,$ we arrive at $R(\mathcal{A},x,y)/\omega(|x-y|)\lesssim\varepsilon,$ which is the desired estimate.
Let us now study the other (non-symmetric) case of (8), that is, when $|y|\geq K,$ and $|x|\leq M.$ Let us write, for every $0\leq k\leq m,$
@@
$\lesssim_{m}\sum_{j=0}^{m-k}\Big{\|}A_{k+j}(y)-\sum_{l=0}^{m-k-j}\frac{1}{l!}A_{k+j+l}(x)(y-x)^{l}\Big{\|}|x-y|^{j}+\max_{1\leq l+j\leq m-k}\|A_{k+j+l}(x)\||x-y|^{l+j}$
$\lesssim_{m}\sum_{j=0}^{m-k}\Big{\|}A_{k+j}(y)-\sum_{l=0}^{m-k-j}\frac{1}{l!}A_{k+j+l}(x)(y-x)^{l}\Big{\|}|x-y|^{j}+\max_{1\leq j\leq m}\|A_{j}(x)\||x-y|^{m-k};$ (12)
where in the last estimate we used that $|x-y|\geq K-M\geq 1$ by virtue of (7), and so $|x-y|^{l+j}\leq|x-y|^{m-k}$ whenever $1\leq l+j\leq m-k$. Now, for the first term in (2), we use the estimates (11) for each $j=0,\ldots,m-k$, but swapping the roles of $x$ and $y$ (note that the estimates of (11) were obtained for $|x|\geq K$ and $|y|\leq M$). We thus obtain,
@@
As for the second term in (2), we use the current assumption $|x|\leq M$ and the second inequality of (7) to infer that
$\max_{1\leq j\leq m}\|A_{j}(x)\||x-y|^{m-k}\leq\max_{1\leq j\leq m}\sup_{u\in E,\,|u|\leq M}\|A_{j}(u)\||x-y|^{m-k}\leq\varepsilon\omega(K)|x-y|^{m-k}.$
But again $|x-y|\geq K-M\geq K/2$ thanks to (7), and so property (3) of $\omega$ tells us that $\omega(K)\lesssim_{\omega}\omega(|x-y|).$
Dividing by $|x-y|^{m-k}\omega(|x-y|)$ in (2) and combining the above two estimates, we conclude that
@@
So, let us assume that $F\in\dot{\operatorname{VC}}^{m,\omega}_{\Gamma}(\mathbb{R}^{n},V)$ and we will next prove that the restriction of $F$ to $E$ satisfies the properties from Definition 1.6. While the proof in the case $\Gamma=\operatorname{small}$ is immediate from Taylor’s theorem, for the scales $\Gamma=\operatorname{large}$ and $\Gamma=\operatorname{far}$, we need to study a couple of subcases separately.
@@
Let us briefly remind why, in the non-trivial case $10\leq k\leq m-$ 1, estimate (13) holds in the setting of normed-valued $C^{m,\omega}$ mappings. Indeed, for each linear and continuous functional $v^{*}\in V^{*},$ with $\|v^{*}\|_{*}=1$ (meaning the dual norm of $V$), we have that $v^{*}\circ F\in C^{m}(\mathbb{R}^{n},\mathbb{R})$ with
$D^{j}(v^{*}\circ F)(x)=v^{*}\circ D^{j}F(x),\quad x\in\mathbb{R}^{m},\,j=1,\ldots,m.$
Thus, using the notation
$\|v^{*}(A)\|=\sup\{\|v^{*}(A(u_{1},\ldots,u_{k}))\,:\,u_{1},\ldots,u_{k}\in\mathbb{S}^{n-1}\},$
for every $k$-linear mapping $A$ from $\mathbb{R}^{n}$ to $V,$ and defining
$A_{x,y,k}:=D^{k}F(x)-\sum_{j=0}^{m-k}\frac{1}{j!}D^{k+j}F(y)(x-y)^{j},\quad x,y\in\mathbb{R}^{n},\,x\neq y,\,k=0,\ldots,m-1,$
Taylor’s theorem applied to $v^{*}\circ F$ gives us, for all $k=0,\ldots,m-1,$
$\frac{\|v^{*}\left(A_{x,y,k}\right)\|}{\omega(|x-y|)|x-y|^{m-k}}=\frac{\Big{\|}D^{k}(v^{*}\circ F)(x)-\sum_{j=0}^{m-k}\frac{1}{j!}D^{k+j}(v^{*}\circ F)(y)(x-y)^{j}\Big{\|}}{\omega(|x-y|)|x-y|^{m-k}}$
$\quad\leq\frac{1}{(m-k)!}\sup_{z\in[x,y]}\frac{\|D^{m}(v^{*}\circ F)(z)-D^{m}(v^{*}\circ F)(y)\|}{\omega(|x-y|)}$
$\quad=\frac{1}{(m-k)!}\sup_{z\in[x,y]}\frac{\|v^{*}\circ D^{m}F(z)-v^{*}\circ D^{m}F(y)\|}{\omega(|x-y|)}\leq\frac{1}{(m-k)!}\sup_{z\in[x,y]}\frac{\|D^{m}F(z)-D^{m}F(y)\|}{\omega(|x-y|)}.$
The Hahn-Banach Theorem in $V$ provided us, for each $k,x,y$ as above, with $v^{*}\in V^{*}$ so that $\|v^{*}\|_{*}=1$ and $\|v^{*}(A_{x,y,k})\|=\|A_{x,y,k}\|.$ Together with the previous estimates, this yields (13).
@@
Also, by the condition (5), we can find $M>0$ (depending on $\varepsilon$, $R$ and $\omega$) so that
$M\geq 2R\quad\text{and}\quad\omega(2R)\leq\varepsilon\omega(M/2).$ (15)
Now, let $x,y\in\mathbb{R}^{n}$ be such that $|x|,|y|\geq M$. By the continuity of $D^{m}F,$ we can find a point $z\in[x,y]$ maximizing the supremum of the right-hand side of (13). In the case where $\max(|y|,|z|)\geq R,$ then